An urn contains 8 white balls and 4 red balls. In how many ways can 4 balls be drawn if
a) 3 balls are white an 1 is red
b) 2 balls are white and 2 are red
c) all four balls are white

Answer by stanbon(75887)

An urn contains 8 white balls and 4 red balls. In how many ways can 4 balls be drawn if
a) 3 balls are white an 1 is red
[8C3][4C1]= 56*4=224
b) 2 balls are white and 2 are red
[8C2][4C2]=28*6=168
c) all four balls are white
8C4=70
--------------
Cheers,
Stan H.